What do misleading graphs contain?
They usually contain a "break" in the graph, which would be on the left side of the graph.
It is 0.3292, approx.
Is course evaluation scale nominal or ordinal or ratio or interval?
It is Ordinal:Order the data from smallest to largest or "worst" to "best".Each data value can be compared with another data value.
What are some uses of statistics in your life?
statistics is used in all aspect of life. so which ever one human finds himself it is alwayz beter to give it a name
examples are agricultural statistics, insurance statistics, actuarial statistics,educational statistics,financial statistics etc
How do you find the meadian of an even amount of numbers?
Assuming that you are referring to the median, the answer is to arrange the number is size-order and take the arithmetic mean (average) of the middle two numbers.
What is the front-end estimation for 2356?
2,000. A front-end estimation just means to round it by the greatest value, in this case the thousands.
Out of the 36 possible and equally-probable outcomes (1st die's number and 2nd die's number), only 3 combinations give you a sum greater than 10:
* 1st die gives a 5 and 2nd die gives a 6
* 1st die gives a 6 and 2nd die gives a 5
* 1st die gives a 6 and 2nd die gives a 6
Therefore the probability that the sum is greater than 10 equals 3/36, or 1/12.
How many codes are there that have 3 letters and 2 numbers?
There are 26 different letters.
There are 10 different numbers.
If a letter or number is allowed to repeat in the code:
Number of ways to make 3 letters = 26 x 26 x 26 = 17,576
Number of ways to make 2 numbers = 10 x 10 = 100
Number of ways to make 3 letters and 2 numbers = 17,576 x 100 = 1,757,600
Each group of 3 letters and 2 numbers can be arranged in (5 x 4 x 3 x 2) = 120 different ways.
So the number of different codes = 1,757,600 x 120 = 210,912,000 .
If a letter or number is not allowed to repeat in the code:
Number of ways to make 3 letters = 26 x 25 x 24 = 15,600
Number of ways to make 2 numbers = 10 x 9 = 90
Number of ways to make 3 letters and 2 numbers = 15,600 x 90 = 1,404,000
Each group of 3 letters and 2 numbers can be arranged in (5 x 4 x 3 x 2) = 120 different ways.
So the number of different codes = 1,404,000 x 120 = 168,480,000 .
What the probability of rolling a sum of nine with a single toss of two fair six sided dice?
There are 36 different ways to roll 2 6-sided dice.
The ones that make 9 are:
3+6 , 4+5 , 5+4 , 6+3
4 out of 36 = 4/36 = 0,111111111 = 11,1 %
What are the advantages and drawbacks of using gold standard?
Unfortunately there are disadvantages to the gold standard. One of the main disadvantages of implementation is that a gold standard would artificially inflate gold's value, increasing the cost of items and industrial process in which it is used. Another disadvantage is under the gold standard, gold mined at a different rate than the economy grows can produce both inflation, when deposits are discovered and extracted and deflation when they are mined to exhaustion
What is the probability of rolling two fair value dice getting two different numbers?
Pr(Two different numbers) = 1 - Pr(Two same) = 1 - 1/6 = 5/6 = 83.3%
How many even 4 digit numbers can you form from the digits 1 2 3 and 5?
In order for the number to be even, the '2' must be the last digit. So we only need to
figure out how many ways there are to arrange the 1, 3, and 5 in the first 3 places.
The first place can be any one of 3 digits. For each of those ...
The second place can be either one of the 2 remaining digits.
And then the third place has to be the only digit that's left.
So there are (3 x 2) = 6 ways to arrange 1, 3, and 5 in the first 3 places.
That means only 6 even 4-digit numbers can be made from the digits 1, 2, 3, and 5.
And here they are:
1 3 5 2
1 5 3 2
3 1 5 2
3 5 1 2
5 1 3 2
5 3 1 2
How many possible outcomes are there of choosing 1 number from the first 6 positive whole numbers?
Six.
What incidence rate and prevalence rate?
prevalencerefers to the number of cases of disease thatexistsin a population and the incidence refers to the number of new cases of diseaseoccurringin a population.
The most important thing that the store owner should learn is that a stem and leaf plot is a huge waste of time in her circumstances. She ought to learn that using some fancy statistical technique may not always be appropriate. There are other methods that are far more useful.
A stem-and-leaf plot will tell her the range of prices of the goods that she sells, along with median and percentiles, etc. but what use is that information to a storekeeper? They may want to know what lines are selling well, which ones are not moving at all. Even more important, they may want to know how the high-profit-margin lines are doing.
Alternatively, she may wish to compare her prices with her competition, but that is useful only if she can compare prices of the same (or similar) goods. A stem-and-leaf plot does not identify the individual product lines and so she would not be able to undertake such a comparison.
All in all, I cannot see a single reason for a stem-and-leaf plot - other than to give her school-age child something to do!
What are the odds of winning solitaire?
playing over 15k games, i have a success ratio of >53% (with undo)
so the real possibilities of winning are even higher!
The Mayflower Compact
What other numbers belong with these numbers 348 6741 and 949?
I want an answer to my question two njasdfklfhs hdjsalfh shut my down
When a coin is tossed why it appears head or tail?
Those are the names given to the two faces of a coin. From early days, coins had an image of important persons - kings, queens, emperors, on one face and that face was called heads. The reason for the other side being called "tails" is uncertain.
These are the only two outcomes of a coin being tossed. Although there is a non-zero probability of the coin falling and remaining on its edge, it is small enough to be ignored in practice.